# MoD-SLAM: Monocular Dense Mapping for Unbounded 3D Scene Reconstruction

### 26 Feb 2024 v4

@@ 4.1 Experimental Setup
Datasets For our quantitative analysis, we implement our system on the Replica dataset [28], ScanNet dataset [7] and TUM RGB-D datase [29]. We follow the same pre-processing step for datasets as in NICE-SLAM [46]. Each dataset presents challenging camera dynamics, effectively showcasing our model’s performance and accuracy. We concurrently utilize these datasets to assess our model in both monocular RGB and RGB-D.
@@
Implementation Details We run our system on a PC with a 4.70 GHz AMD 7900X CPU and NVIDIA RTX 4090 GPU. We use the pre-trained model from DROID-SLAM [34] for tracking and use the same parameters in the tracking process as it. In the reconstruction process, we choose to sample a ray at several points of $N_{strat}=20$ and $N_{imp}=40$. For loss weights, we choose $\lambda_{c}=1.0$, $\lambda_{dep}=1.0$, $\lambda_{dist}=0.002$, $\lambda_{eik}=0.15$, $\lambda_{sdf}=1.0$. The compared baseline includes classical SLAM systems ORB-SLAM3 [4] and DROID-SLAM [30], newest SLAM systems NICE-SLAM [46], HI-SLAM [42] and GO-SLAM [43] and 3D Gaussian-based SLAM Photo-SLAM [11].
@@
NICER-SLAM[45] 1.88 3.65 25.41 0.827 6
Mono GO-SLAM[43] 0.39 4.39 3.81 22.13 0.733 8
Photo-SLAM[11] 1.09 33.30 0.930 42
Ours 0.35 3.23 2.48 28.37 0.850 8
@@
GO-SLAM[43] 5.35 7.52 7.03 7.74 6.84 5.26 4.78 6.36
RGB-D Co-SLAM[38] 7.13 11.14 9.36 5.90 11.81 7.14 6.77 8.47
E-SLAM[12] 7.32 8.57 7.53 6.43 8.85 5.74 5.13 7.08
Ours 5.27 7.44 6.73 6.48 6.14 5.31 3.29 5.81
Table 2: ATE RMSE [cm] on the ScanNet dataset [7]. The performance metrics reported by our system for each scene are derived from the average of three executions.
@@
Scene fr1-desk fr2-xyz fr3-office
GO-SLAM[43] 1.500 0.600 1.000
RGB-D Co-SLAM[38] 2.400 1.700 2.400
Ours 1.382 0.641 0.939
DROID-SLAM[34] 2.713 1.304 2.573
DeepV2D[33] 21.391 6.950 11.922
Mono ORB-SLAM3[4] 1.534 0.720 1.400
GO-SLAM[43] 2.381 1.093 1.773
Ours 1.481 0.686 1.101
@@ 4.2 Results and Evaluation
Evaluation on ScanNet [7] We evaluated our system across seven distinct scenes in ScanNet dataset [7]. Table 2 presents the tracking results. Our system outperforms others in tracking accuracy, particularly in the large indoor scenes of ScanNet. As illustrated in Fig. 6 and Fig. 5, our system outmatches GO-SLAM [43] by rendering the scene with enhanced detail fidelity, more accurately approximating the real-world geometry and ensuring smoother scene transitions. Furthermore, our framework possesses a pronounced proficiency in addressing scene sparsity through effective void interpolation.
Evaluation on TUM-RGBD [29] To further test our method, we compared our method with point-cloud SLAM methods [34, 33, 4] and NeRF-based SLAM methods [43, 38] on the TUM RGB-D dataset. Since the TUM RGB-D dataset lacks ground truth meshes, we compared the accuracy of the tracking process. Table 3 shows the accurate tracking performance of our method.
Runtime Analysis We tested the tracking and the mapping time of NICER-SLAM [45], GO-SLAM [43] and our system using the same number of sampling points. Fig. 7(b) shows that our system is faster than GO-SLAM and NICER-SLAM in the tracking process. However, because of the depth estimation module, our system is slightly slower than GO-SLAM in the mapping process, but higher than NICER-SLAM. In addition, Fig. 7(b) also demonstrates that our model size is smaller and uses less memory.
@@ 5 Conclusion
We presented MoD-SLAM, a monocular implicit neural-based dense mapping method that demonstrates state-of-the-art accuracy in monocular reconstruction. Our experiments show that by integrating a monocular depth estimation module, our system not only achieves a more accurate scene-scale reconstruction but also excels in reconstructing the details and textures of geometric objects compared to previous SLAM systems. Additionally, we show the superior performance of our system in reconstructing unbounded scenes after employing Gaussian encoding and reparametrization techniques.
Limitations The depth estimation module relies on a pre-trained model; thus, for complex scenes or those significantly divergent from our provided pre-trained scenes, it’s necessary to retrain the depth estimation module. Moreover, compared to traditional SLAM, our method lacks an IMU [32] module, leading to tracking errors.
@@ Appendix 0.A Appendix
0.A.1 Ablation Study
In this section, we evaluate the choice of our depth estimation module, multivariate Gaussian encoding and reparameterization.
@@ 0.A.1 Ablation Study
Unbounded Scene We also conduct tests to assess the impact of reparameterization Equ. 6 and Gaussian encoding Equ. 5 on reconstructing scenes with no defined boundaries. For comparison, we used NEWTON [14] as a baseline to evaluate our system’s performance. The quantitative data in Table 6 indicates that our reparameterization and Gaussian encoding significantly improve the accuracy in both scaling and geometric reconstruction of unbounded scenes. These approaches notably enhance our system’s overall performance. Our data were acquired using the RGB-D mode of our system within the ScanNet 0207 scene. NEWTON’s data are from the original paper [14]. The results on the ScanNet [7] indicate that our model shows greater robustness and more precise mapping abilities compared to NEWTON.
@@
1 1 0 0 0 85.79 27.76
1 1 1 0 0 85.98 27.89
1 1 1 1 0 89.01 28.13
1 1 1 1 1 89.73 28.37
Table 5: Depth ablation study. We explore the effects of the depth estimation module (DepE.) and the depth distillation module (DepD.) in our system.
@@
None - -
NEWTON[14] 23.15 0.702
Table 5: Depth ablation study. We explore the effects of the depth estimation module (DepE.) and the depth distillation module (DepD.) in our system.
@@
0.A.2 Additional Experimental Results
Replica Dataset [29] Here we show more results on Replica [29]. Fig. 8, Fig. 9 and demonstrate the superior scene reconstruction capability of our proposed method in monocular and RGB-D mode, and also exhibiting enhanced smoothness in comparison to GO-SLAM [43].

### 08 Mar 2024 v5

@@ 4.1 Experimental Setup
Datasets For our quantitative analysis, we implement our system on the Replica dataset [28] and ScanNet dataset [7]. We follow the same pre-processing step for datasets as in NICE-SLAM [45]. Each dataset presents challenging camera dynamics, effectively showcasing our model’s performance and accuracy. We concurrently utilize these datasets to assess our model in both monocular RGB and RGB-D.
@@
Implementation Details We run our system on a PC with a 4.70 GHz AMD 7900X CPU and NVIDIA RTX 4090 GPU. We use the pre-trained model from DROID-SLAM [34] for tracking and use the same parameters in the tracking process as it. In the reconstruction process, we choose to sample a ray at several points of $N_{strat}=20$ and $N_{imp}=40$. For loss weights, we choose $\lambda_{c}=1.0$, $\lambda_{dep}=1.0$, $\lambda_{dist}=0.002$, $\lambda_{eik}=0.15$, $\lambda_{sdf}=1.0$. The compared baseline includes classical SLAM systems [4, 30], newest SLAM systems [45, 42, 44, 37, 12] and 3D Gaussian-based SLAM system [11].
@@
[…]
Table 2: ATE RMSE [cm] on ScanNet [7]. The performance metrics reported by our system for each scene are derived from the average of three executions.
@@
[…]
@@ 4.2 Results and Evaluation
Evaluation on ScanNet [7] We evaluated our system across seven distinct scenes in ScanNet dataset [7]. Table 2 presents the tracking results. Our system outperforms others in tracking accuracy, particularly in the large indoor scenes of ScanNet. As illustrated in Fig. 6 and Fig. 5, our system outmatches GO-SLAM [42] by rendering the scene with enhanced detail fidelity, more accurately approximating the real-world geometry and ensuring smoother scene transitions. Furthermore, our system possesses a pronounced proficiency in addressing scene sparsity through effective void interpolation.
Runtime Analysis We test the tracking and the mapping time of NICER-SLAM [44], GO-SLAM [42] and our system using the same number of sampling points. Fig. 7(b) shows that our system is faster than GO-SLAM and NICER-SLAM in the tracking process. However, because of the depth estimation module, our system is slightly slower than GO-SLAM in the mapping process, but higher than NICER-SLAM. In addition, Fig. 7(b) also demonstrates that our model size is smaller and uses less GPU memory.
@@ 5 Conclusion
We presented MoD-SLAM, a monocular unbounded NeRF-based dense mapping method that demonstrates state-of-the-art performance. The proposed Gaussian-based unbounded scene representation approach allows our system to be capable of reconstructing unbounded scenes. The depth estimation module and depth distillation module provide prior depth values, addressing scale inconsistency in monocular unbounded scene reconstruction. The depth-supervised camera tracking method combines traditional SLAM method with the depth loss term, which achieves robust and accurate camera tracking and pose estimation. Our experiments show that MoD-SLAM has superior performance in both localization and reconstruction with low time and GPU memory consumption compared to state-of-the-art SLAM systems.
@@ Appendix 0.A Appendix
0.A.1 Loop Closure Detection
0.A.2 Ablation Study
In this section, we evaluate the choice of our depth estimation module, multivariate Gaussian encoding, depth loss term, and reparameterization.
@@ 0.A.2 Ablation Study
Unbounded Scene We also conduct tests to assess the impact of reparameterization Equ. 4 and Gaussian encoding Equ. 3 on reconstructing scenes with no defined boundaries. For comparison, we used NEWTON [14] as a baseline to evaluate our system’s performance. The quantitative data in Table 5 indicates that our reparameterization and Gaussian encoding significantly improve the accuracy in both scaling and geometric reconstruction of unbounded scenes. These approaches notably enhance our system’s overall performance. Our data were acquired using the RGB-D mode of our system within the ScanNet 0207 scene. NEWTON’s data are from the original paper [14]. The results on the ScanNet [7] indicate that our model shows greater robustness and more precise mapping abilities compared to NEWTON.
Camera Tracking We explore the effect of the depth loss term on camera localization. To verify the robustness of our proposed depth loss term, we further explore the results of solely applying L1 loss or L2 loss on the system. Table 6 shows the experimental results on the Replica dataset [28]. Our proposed depth loss term can significantly improve the camera tracking accuracy of the system. The convergence performance of L1 loss is poor, making it unsuitable for real-time SLAM systems. L2 loss is overly sensitive to depth outliers, hence not suitable for our depth estimation module-based SLAM system.
@@
[…]
Table 4: Depth modules ablation study. We explore the effects of the depth estimation module (DepE.) and the depth distillation module (DepD.) in our system.
@@
Table 6: Camera Tracking Ablation Study. We report ATE RMSE[cm] on the Replica [28]. Experimental results demonstrate that the loss function proposed by us exhibits superior robustness.
office0 office1 office2 office3 office4 room0 room1 room2 Avg.
None 0.310 0.327 0.339 0.512 0.603 0.588 0.391 0.307 0.422
[…]
Scene fr1-desk fr2-xyz fr3-office
GO-SLAM[42] 1.500 0.600 1.000
[…]
DROID-SLAM[34] 2.713 1.304 2.573
[…]
0.A.3 Additional Experimental Results
Evaluation on TUM-RGBD [29] To test our method further, we compared our method with point-cloud SLAM methods [34, 33, 4] and NeRF-based SLAM methods [42, 37] on the TUM RGB-D dataset. Since the TUM RGB-D dataset lacks ground truth meshes, we compared the accuracy of the tracking process. Table 7 shows the accurate tracking performance of our method.
Evaluation on Replica [28] Here we show more results on Replica [28]. Fig. 8, Fig. 9 and demonstrate the superior scene reconstruction capability of our proposed method in monocular and RGB-D mode, and also exhibiting enhanced smoothness in comparison to GO-SLAM [42].